# Loss Gradient Gaussian Width based Generalization and Optimization Guarantees

### v1

@@ 2.2 Main Result: Gaussian Width Bound
$\sup_{\hat{\xi}^{(n)}\in\hat{\Xi}^{(n)}}\frac{1}{\sqrt{n}}\left\|\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}\right\|_{2}\leq c_{5}(1+u)w(\hat{\Xi}_{n})~{}.\vspace*{-2mm}$ (9)
@@
${\cal L}_{{\cal D}}(\hat{\theta})-{\cal L}_{{\cal D}}(\theta^{*})\leq 2\bar{c}_{1}\left(\|\nabla\hat{{\cal L}}_{n}(\hat{\theta})\|_{2}+2\left(\frac{4c_{6}w(\hat{\Xi}_{n})}{\sqrt{n}}+c_{3}\frac{\log\frac{1}{\delta}}{n}\right)\right)+c_{4}\sqrt{\frac{\log\frac{1}{\delta}}{n}}~{}.$ (11)
@@
A technically curious aspect of the result in (10) is that the randomness in $\hat{R}_{n}(\hat{\Xi}^{(n)})$ is from $n$ Rademacher variables over samples whereas that in $w(\hat{\Xi}_{n})$ is from $p$ normal variables over gradient components. The shift from samples to gradient components makes the bound geometric. ∎
@@
Remark 2.4.
For the result to be meaningful, $\Theta$ needs to be suitably chosen, without using the training set. For example, for deep learning models, $\Theta$ can constitute parameters encountered in the last 10% of epochs during training. There is empirical evidence showing loss gradients close to convergence have small norms and the loss Hessians, which capture the change in gradients, have sharply decaying eigenvalues. We review and build on such results in Section 4. ∎
@@ 2.3 Proof sketch of Main Result
Our analysis is based on generic chaining (GC) and needs the following key definition due to [80].
@@
  • For the set $\hat{\Xi}_{n}$ of average gradients $\hat{\xi}_{n}$, we construct a Gaussian process $X_{\hat{\xi}_{n}}=\langle\hat{\xi}_{n},\mathbf{g}\rangle$ where $\mathbf{g}\sim{\cal N}(0,\mathbb{I}_{p\times p})$. We note a simple one-to-one correspondence between the set of average gradients $\hat{\Xi}_{n}$ and the set of $n$-tuples $\hat{\Xi}^{(n)}$ since $\hat{\xi}_{n}=\frac{1}{n}\sum_{i=1}^{n}\xi_{i}$. Further, we show that the canonical distance $d(X_{\hat{\xi}_{1,n}},X_{\hat{\xi}_{2,n}})$ as in (14) from the Gaussian process on $\hat{\Xi}_{n}$ satisfies $d(X_{\hat{\xi}_{1,n}},X_{\hat{\xi}_{2,n}})=\bar{d}^{(n)}(\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2})$, the non-canonical distance on $\hat{\Xi}^{(n)}$. Based on the set correspondence and exact same distance measure, we have $\gamma_{2}(\hat{\Xi}_{n},d)=\gamma_{2}(\hat{\Xi}^{(n)},\bar{d}^{(n)})$, so that $\hat{R}_{n}(\hat{\Xi}^{(n)})\leq c_{2}\frac{\gamma_{2}(\hat{\Xi}_{n},d)}{\sqrt{n}}$.
@@
We provide additional details and associated Lemma for the first part of the analysis. Consider the stochastic process $X_{\hat{\xi}^{(n)}}:=\frac{1}{\sqrt{n}}\left\|\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}\right\|_{2}$, by triangle inequality, we have
@@
which is greater or equal to the canonical distance as in (14), due to Jensen’s inequality in (15). With our choice of distance $\bar{d}^{(n)}$, note that $\mathbf{v}_{i}=\frac{\xi_{i,1}-\xi_{i,2}}{\bar{d}^{(n)}(\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2})}$ satisfies $\sum_{i=1}^{n}\|\mathbf{v}_{i}\|^{2}_{2}=n$. Then, the following result is effectively a shifted increment condition (SIC), a shifted variant of (16) with a constant shift of $\mu\leq 1$:
@@
$\mathbb{P}_{\varepsilon^{(n)}}\left(\left|\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}\mathbf{v}_{i}\right\|_{2}-\mu\right|\geq u\right)\leq 2\exp\left(-\frac{u^{2}}{2}\right),$ (18)
@@
The $X_{\hat{\xi}^{(n)}_{0}}=0$ is easy to satisfy for gradients as long as the gradients are all zero for some $\theta_{0}$, e.g., minima of the loss, interpolation condition, etc. That condition is not necessary, and a more general result can be established, e.g., see [80][Theorem 2.4.12]. ∎
@@
The SIC in Lemma 1 and the shifted variant of GC in Lemma 20 can now be used to establish the first part of the proof of Theorem 1. The technicalities for the other parts are more straightforward and are directly handled in the proof of Theorem 1.
3 SGD with Sample Reuse
@@
Stochastic gradient descent (SGD) updates parameters following $\theta_{t+1}=\theta_{t}-\eta_{t}\nabla\ell(\theta_{t},z_{t}^{(b)})$, where $z_{t}^{(b)}$ are a mini-batch of $b$ fresh samples drawn from the underlying distribution ${\cal D}$. The mini-batch stochastic gradient is assumed to be unbiased. In practice, however, one reuses samples, by repeatedly drawing $z_{t}^{(b)}$ from a fixed set of $n$-samples $z^{(n)}=\{z_{1},\ldots,z_{n}\}$ drawn i.i.d. from ${\cal D}$ at the beginning. Such sample reuse violates the fresh sample assumption, and in turn cannot assume the mini-batch SGs to be unbiased [18, 77]. The theory and practice of SGD and variants have diverged: the theory continues to assume fresh samples [90, 28] whereas in practice one does multiple passes (epochs) over the fixed $z^{(n)}$ [34, 72]. Such sample reuse can be viewed as empirical risk minimization (ERM) [75, 73, 28, 63] where the analysis is focused only on the finite-sum objective, and one needs to use uniform convergence (UC) to get population results [75, 73]. While the machinery of ERM followed by UC is quite well established, the classical UC analysis is usually over large function class, and so cannot quite take advantage of the geometry of individual stochastic gradients [59, 3, 20].
In this section, we provide a generalization and optimization analysis of SGD with sample reuse. We start with the simple case, i.e., gradient descent (GD), and show the deviation of the sample estimate from the population gradient can be bounded in terms of LGGW. We present detailed results and analyses of general mini-batch SGD in Appendix B.2.
Gradient descent. For $t\in[T]$, GD uses the full set $z^{(n)}$ and $\theta_{t+1}=\theta_{t}-\eta_{t}\nabla\ell(\theta_{t},z^{(n)})$, where $\nabla\ell(\theta_{t};z^{(n)})=\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})$. At each epoch $t$, our goal is to characterize the deviation of the sample estimate $\nabla\ell(\theta_{t};z^{(n)})$ from the population gradient $\nabla{\cal L}_{{\cal D}}(\theta_{t})$.
@@
Let $\theta_{t}\in\Theta,t\in[T]$ be a sequence of parameters obtained from GD by reusing a fixed set of samples $z^{(n)}\sim{\cal D}^{n}$ in each epoch. With $w(\hat{\Xi}_{n})$ denoting the LGGW as in Definition 2, for any $\delta\leq\frac{1}{2}$, with probability at least $(1-2\delta)$ over $z^{(n)}\sim{\cal D}^{n}$, we have
$\max_{t\in[T]}\sup_{\theta_{t}\in\Theta}\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta_{t},z_{i})-\nabla{\cal L}_{D}(\theta_{t})\right\|_{2}\leq\frac{c_{7}w(\hat{\Xi}_{n})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{n}}~{},$ (21)
where the population gradient $\nabla{\cal L}_{D}(\theta_{t})=\mathbb{E}_{z\sim{\cal D}}[\nabla\ell(\theta_{t};z)],t\in[T]$.
Note that the result in Theorem 2 holds for each $\theta_{t}$, where $\theta_{t}$ for any $t>1$ is adaptively generated based on the reused samples $z^{(n)}$. Further, our proof does not utilize the specific form of the GD/SGD updates, and in fact goes through for any update of the form $\theta_{t+1}=f(\{\theta_{s},s\leq t\},z^{(b)})$. Thus, Theorem 2 is an adaptive data analysis (ADA) result, but without using the traditional tools such as differential privacy [23, 41]. Our proof carefully utilizes the conditional probability of events, conditioned on suitable past events, and expresses adaptive events in terms of events based on fresh samples. Proofs are in Appendix B. The mini-batch SGD with batch size $b$ will have a similar bound with sample dependence $\frac{1}{\sqrt{b}}$ (Theorem 7 in Appendix B.2), which highlights a computational and statistical trade-off between GD and SGD: GD has a $\frac{1}{\sqrt{n}}$ statistical rate but takes $O(n)$ computation per step whereas SGD has a $\frac{1}{\sqrt{b}}$ statistical rate but only needs $O(b)$ computation per step.
@@
There are a few key takeaways from Theorems 2 and 7. First, note that the key term in the bound is essentially the same in both results: the LGGW $w(\hat{\Xi}_{n})$. As long as the LGGW $w(\hat{\Xi}_{n})$ is small, the bound stays small. Second, the bounds have a benign dependence on the number of steps, in particular $\log T$. Third, the sample dependence is $\frac{1}{\sqrt{n}}$ for GD with batch size $n$ and $\frac{1}{\sqrt{b}}$ for mini-batch SGD with batch size $b$. Larger mini-batch sizes yield a sharper bound. ∎
Population Convergence in Optimization. The results in Theorems 2 can be readily used to straightforwardly go from uniform convergence of empirical gradients during optimization to convergence of population gradients, without needing an additional uniform convergence argument [75, 74].
@@
Consider a non-convex loss ${\cal L}_{{\cal D}}(\theta)=\mathbb{E}_{z\sim{\cal D}}\left[\ell(\theta,z)\right]$ with $\tau$-Lipschitz gradient. For GD using a total of $z^{(n)}\sim{\cal D}^{n}$ original samples and re-using $n$ samples in each step as in Theorem 2, with step-size $\eta=\frac{1}{4\tau}$, with probability at least $(1-2\delta)$ for any $\delta>0$ over $z^{(n)}\sim{\cal D}^{n}$,
$\mathbb{E}_{R}\left\|\nabla{\cal L}_{{\cal D}}\left(\mathbf{\theta}_{R}\right)\right\|^{2}\leq O\left(\frac{1}{T}\right)+O\left(\frac{c_{1}w(\hat{\Xi}_{n})+\sqrt{(\log T+\log\frac{1}{\delta})/2}}{\sqrt{b}}\right),$
@@
The bound has two terms: the first term comes from the finite-sum optimization error, which depends on the number of iterations $T$, and the second term comes from the statistical error, which depends on the sample size $n$ and the LGGW $w(\hat{\Xi}_{n})$. Similar results can be developed under additional assumptions on the loss, e.g., for a $\alpha$-strongly convex with Lipshitz gradient, we can get a similar bound with the first term improving to $\exp(-T/\alpha)$. ∎
@@
While such an analysis for SGD can be done using differential privacy based adaptive data analysis (ADA) [24, 41] and algorithmic stability [16, 36], the best-known efficient algorithm in such framework [12] have gradient error scaling with $\sqrt{p}$. Further, without any assumption, lower bounds in ADA [78] also show that it is hard to release more than $\tilde{O}(n^{2})$ statistical queries. Therefore, capturing the structure in the gradients to obtain sharper bounds for sample reuse is novel and avoids the issues with existing analysis. ∎
4 Gaussian Width of Deep Learning Gradients

### v2

@@ 2.2 Main Result: Gaussian Width Bound
$\sup_{\hat{\xi}^{(n)}\in\hat{\Xi}^{(n)}}\frac{1}{\sqrt{n}}\left\|\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}\right\|_{2}\leq c_{5}(1+u)w(\hat{\Xi}_{n})~{}.$ (9)
@@
$\begin{split}&{\cal L}_{{\cal D}}(\hat{\theta})-{\cal L}_{{\cal D}}(\theta^{*})\leq 2\bar{c}_{1}\left(\|\nabla\hat{{\cal L}}_{n}(\hat{\theta})\|_{2}+2\left(\frac{4c_{6}w(\hat{\Xi}_{n})}{\sqrt{n}}+c_{3}\frac{\log\frac{1}{\delta}}{n}\right)\right)+c_{4}\sqrt{\frac{\log\frac{1}{\delta}}{n}}~{}.\end{split}$ (11)
@@
A technically curious aspect of the result in (10) is that the expectation in $\hat{R}_{n}(\hat{\Xi}^{(n)})$ is from $n$ Rademacher variables over samples whereas that in $w(\hat{\Xi}_{n})$ is from $p$ normal variables over gradient components. The shift from samples to gradient components makes the bound geometric. We elaborate on such geometry in the next remark. ∎
Remark 2.4.
Since the bound is with respect to the Gaussian width of the set of scaled stacked gradients $\hat{\xi}_{n}=\frac{1}{\sqrt{n}}\big{[}\nabla\ell(\theta;z_{i})\big{]}_{i=1}^{n}$, it would depend on both the alignment of the $n$ sample gradients $\xi_{i}=\nabla\ell(\theta;z_{i})$ as well as the width of the set of individual sample gradients $\Xi_{i}=\left\{\hat{\xi}_{i}\in\mathbb{R}^{p}\mid\exists\theta\in\Theta,~{}\text{s.t.}~{}~{}\hat{\xi}_{i}=\nabla\ell\left(\theta;z_{i}\right)\right\}$. Depending on the alignment over the $n$ sample gradients, the width can range from $O(1)$ to $O(\sqrt{n})$, similar to Rademacher complexity [74, 10]. The nature of alignment here will be based on similarity of the hierarchical covering of individual gradient sets, with similar coverings, e.g., similar distortions at each level of the hierarchy, yielding small width. For the width of individual sample gradient sets, we present bounds for two commonly used networks: FeedForward Networks and Residual Networks in Section 4. ∎
@@ 2.3 Proof sketch of Main Result
Our analysis is based on generic chaining (GC) and needs the following definition [81]:
@@
  • For the set $\hat{\Xi}_{n}$ of stacked gradients $\hat{\xi}_{n}$, we construct a Gaussian process $X_{\hat{\xi}_{n}}=\langle\hat{\xi}_{n},\mathbf{g}\rangle$ where $\mathbf{g}\sim{\cal N}(0,\mathbb{I}_{np\times np})$. We note a simple one-to-one correspondence between the set of stacked gradients $\hat{\Xi}_{n}$ and the set of $n$-tuples $\hat{\Xi}^{(n)}$ since $\hat{\xi}_{n}=\frac{1}{\sqrt{n}}\left[\xi_{i}\right]_{i=1}^{n}$. Further, we show that the canonical distance $d(X_{\hat{\xi}_{1,n}},X_{\hat{\xi}_{2,n}})$ as in (14) from the Gaussian process on $\hat{\Xi}_{n}$ satisfies $d(X_{\hat{\xi}_{1,n}},X_{\hat{\xi}_{2,n}})=\bar{d}^{(n)}(\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2})$, the non-canonical distance on $\hat{\Xi}^{(n)}$. Based on the set correspondence and exact same distance measure, we have $\gamma_{2}(\hat{\Xi}_{n},d)=\gamma_{2}(\hat{\Xi}^{(n)},\bar{d}^{(n)})$, so that $\hat{R}_{n}(\hat{\Xi}^{(n)})\leq c_{2}\frac{\gamma_{2}(\hat{\Xi}_{n},d)}{\sqrt{n}}$.
@@
We provide additional details and associated lemmas for the first part of the analysis. Consider the stochastic process $X_{\hat{\xi}^{(n)}}:=\frac{1}{\sqrt{n}}\left\|\sum_{i=1}^{n}\varepsilon_{i}\xi_{i}\right\|_{2}$, by triangle inequality, we have
@@
which is greater or equal to the canonical distance as in (14), due to triangle inequality in (15). With our choice of distance $\bar{d}^{(n)}$, note that $\mathbf{v}_{i}=\frac{\xi_{i,1}-\xi_{i,2}}{\bar{d}^{(n)}(\hat{\xi}^{(n)}_{1},\hat{\xi}^{(n)}_{2})}$ satisfies $\sum_{i=1}^{n}\|\mathbf{v}_{i}\|^{2}_{2}=n$. Then, the following result is effectively a shifted increment condition (SIC), a shifted variant of (16) with a constant shift of $\mu\leq 1$:
@@
$\mathbb{P}_{\varepsilon^{(n)}}\left(\left|\left\|\frac{1}{\sqrt{n}}\sum_{i=1}^{n}\varepsilon_{i}\mathbf{v}_{i}\right\|_{2}-\mu\right|\geq u\right)\leq 2\exp\left(-\frac{u^{2}}{2}\right)~{},$ (18)
@@
The $X_{\hat{\xi}^{(n)}_{0}}=0$ is easy to satisfy as long as the gradients are all zero for some $\theta_{0}$, e.g., minima of the loss, interpolation condition, etc. That condition is not necessary, and a more general result can be established, e.g., see [81][Theorem 2.4.12]. ∎
@@
The SIC in Lemma 1 and the shifted variant of GC in Lemma 20 can now be used to establish the first part of the proof of Theorem 1. The technicalities for the other parts are more straightforward and are directly handled in the proof of Theorem 1. All proofs are in Appendix A.
3 Gradient Descent with Sample Reuse
@@
Gradient descent (GD) updates parameters following $\theta_{t+1}=\theta_{t}-\eta_{t}\nabla\hat{{\cal L}}_{n}(\theta_{t})$. If the goal is to make sure GD decreases the population loss ${\cal L}(\theta)$, then one typically assumes that fresh samples are used in every iteration $t$ to compute ^$\hat{{\cal L}}_{n}(\theta_{t})$ and corresponding ^$\nabla\hat{{\cal L}}_{n}(\theta_{t})$, making sure such estimates are unbiased estimates of the population loss gradient. In practice, however, one reuses samples from the fixed training set for GD and also for mini-batch stochastic gradient descent. Such sample reuse violates the fresh sample assumption, and in turn cannot guarantee the unbiasedness of the gradient estimates [19, 78]. While one can still guarantee decrease of the finite sum empirical loss from the ERM perspective [76, 29, 64] with GD, it is unclear if the population loss should decrease with such updates. Thus, the theory and practice of (stochastic) GD and variants have diverged, since the theory continues to assume fresh samples [91, 29] whereas in practice one does multiple passes (epochs) over the fixed dataset [35, 73]. If ERM optimization is done via sample reuse, one needs to then use uniform convergence (UC) to get population results [76, 74]. While this approach is well established, UC analysis over large function classes cannot quite take advantage of the geometry of gradients [60, 3, 21].
In this section, we show that empirical estimates of the gradient with sample reuse stay close to the population gradients as long as the LGGW is small. The consequence of sample reuse is that $\theta_{t}$ was determined using the samples $\{z_{i},i\in[n]\}$ in prior iterations and the gradient estimate ^$\nabla\hat{{\cal L}}_{n}(\theta_{t})=\frac{1}{n}\sum_{i=1}^{n}\ell(\theta_{t},z_{i})$ also uses the same samples. Since $\theta_{t}$ is not independent of the samples, basic concentration bounds such as Hoeffding or McDiarmid are not directly applicable to ^$\hat{{\cal L}}_{n}(\theta_{t})$. The sample reuse here is a specific case of adaptive data analysis, and has been acknowledged and studied in the adaptive data analysis literature [25, 79].
We start by consider the joint event of interest under such sequential sample reuse. The initialization parameter $\theta_{0}$ for gradient descent is assumed to be independent of the samples, e.g., based on random initialization or a pre-trained model. But such a parameter can land anywhere in some $\Theta$, so we need a uniform bound on that set. The subsequent parameters $\theta_{t}=\theta_{t-1}-\eta\nabla\hat{{\cal L}}_{n}(\theta_{t-1})$ are obtained sequentially and thus has dependency on all the samples, which will be reused to estimate ^$\nabla\hat{{\cal L}}_{n}(\theta_{t})$. Our main result shows that the deviation between the empirical gradient and the population gradient jointly over the $(T+1)$ time steps can be bounded with high probability based on the LGGW and otherwise logarithmic dependencies on the ambient dimensionality $p$ and the time horizon $T$.
@@
Let $\theta_{0}\in\Theta_{0}$, and $\theta_{t},t\in[T]$ be a sequence of parameters obtained from GD by reusing a fixed set of samples $z^{(n)}\sim{\cal D}^{n}$ in each epoch. Let
$\Delta(\theta):=\left\|\frac{1}{n}\sum_{i=1}^{n}\nabla\ell(\theta,z_{i})-\nabla{\cal L}_{D}(\theta)\right\|_{2}~{},$ (21)
where the population gradient $\nabla{\cal L}_{D}(\theta_{t})=\mathbb{E}_{z\sim{\cal D}}[\nabla\ell(\theta_{t};z)],t\in[T]$. Under Assumption 2, with $w(\hat{\Xi}^{0}_{n})$ denoting the LGGW for $\Theta_{0}$ as in Definition 2, for any $\delta\leq\frac{1}{2}$, with probability at least $(1-2\delta)$ over $z^{(n)}\sim{\cal D}^{n}$, we have
$\begin{split}\max&\left(\sup_{\theta_{0}\in\Theta_{0}}\Delta(\theta_{0}),~{}~{}\max_{t\in[T]}\Delta(\theta_{t})\right)\leq\frac{c_{7}\max(w(\hat{\Xi}^{0}_{n}),\log p)+\sqrt{\log T+\log\frac{1}{\delta}}}{\sqrt{n}}~{}.\end{split}$ (22)
Our proof does not utilize the specific form of GD updates, and in fact goes through for any update of the form $\theta_{t+1}=f(\{\theta_{s},s\leq t\})$. Thus, Theorem 2 is an adaptive data analysis (ADA) result, but without using the traditional tools such as differential privacy [24, 42]. Our proof utilizes the conditional probability of events, and expresses conditional or adaptive events in terms of events based on fresh samples. Proofs are in Appendix B.
@@
There are a few key takeaways from Theorem 2. First, note that the key term in the bound depends on LGGW $w(\hat{\Xi}^{0}_{n})$ only at initialization, since we considered a uniform bound over $\theta_{0}\in\Theta_{0}$. Note that $w(\hat{\Xi}^{0}_{n})$ can be bounded based on properties of gradients at random initialization or with pretrained models (see Section 4). Second, the bound has a logarithmic dependence on $p$, the ambient dimension, and $T$, the number of steps. Third, the sample dependence is $\frac{1}{\sqrt{n}}$ for GD with batch size $n$. Thus, the bound says that in spite of sample reuse, the sample gradients will stay close to population gradient. ∎
Remark 3.2.
The bound in Theorem 2 does not restrict the subsequent iterates $\theta_{t}$ to be restricted to the initialization set $\Theta_{0}$. In other words, the bound holds without a “near initialization” type restricted which has been widely used in the deep learning theory literature. Further, in case the analysis starts from a given fixed $\theta_{0}$, the dependence on $w(\hat{\Xi}^{0}_{n})$ drops out, and we are only left with dependence on $(\log p+\sqrt{\log T})$. ∎
Population Convergence in Optimization. The results in Theorems 2 can be readily used to get optimization convergence of population gradients, without needing an additional uniform convergence argument [74].
@@
Consider a non-convex loss ${\cal L}_{{\cal D}}(\theta)=\mathbb{E}_{z\sim{\cal D}}\left[\ell(\theta,z)\right]$ with $\tau$-Lipschitz gradient. For GD using a total of $z^{(n)}\sim{\cal D}^{n}$ original samples and re-using the $n$ samples in each step as in Theorem 2, with step-size $\eta=\frac{1}{4\tau}$, with probability at least $(1-2\delta)$ for any $\delta>0$ over $z^{(n)}\sim{\cal D}^{n}$,
$\mathbb{E}_{R}\left\|\nabla{\cal L}_{{\cal D}}\left(\mathbf{\theta}_{R}\right)\right\|^{2}\leq O\left(\frac{1}{T}\right)+O\left(\frac{\max\left(w^{2}(\hat{\Xi}^{0}_{n}),\log^{2}p\right)+\left(\log T+\log\frac{1}{\delta}\right)}{n}\right),$
@@
In the bound, the first term comes from the finite-sum optimization error, which depends on the number of iterations $T$; and the second term comes from the statistical error, which depends on the sample size $n$ and the LGGW $w(\hat{\Xi}^{0}_{n}),\log p$, and $\log T$. ∎
@@
While an analysis of sample reuse based GD can be done using differential privacy (DP) based adaptive data analysis (ADA) [25, 42] and algorithmic stability [17, 37], the best-known efficient algorithm in such framework [13] have gradient error scaling with $\sqrt{p}$. Further, without any assumption, lower bounds in ADA [79] also show that it is hard to release more than $\tilde{O}(n^{2})$ statistical queries. Our analysis of sample reuse without using DP for the special case of estimating high-dimensional gradients leads to sharper bounds, with only logarithmic dependence on the ambient dimensionality. ∎
4 Gaussian Width of Deep Learning Gradients for Neural Networks